OpenStudy (anonymous):

please help me solve for x: 10(1-x/10)-(10-x)-1/100(10-x)=0.05

jhonyy9 (jhonyy9):

10(1-x/10)-(10-x)-1/100(10-x)=0,05 10(1-x) 1 ------ -(10-x)- ------- (10-x) =0,05 you have thought it hence ? 10 100

OpenStudy (anonymous):

I managed to break it down to 10-x-10+x-1/10+x=1/20 (dunno if that is correct) then I combined like terms like I have bee taught.

jhonyy9 (jhonyy9):

you have wrote it very mixed first write me please how is correct this exercise ?

OpenStudy (anonymous):

I dont understand your response

jhonyy9 (jhonyy9):

write me your exersice correct with parantheses

OpenStudy (anonymous):

The original equation or what I have managed to do?

jhonyy9 (jhonyy9):

yes the original equation

jhonyy9 (jhonyy9):

how is correct ?

OpenStudy (anonymous):

10(1-x/10)-(10-x)-1/100(10-x)=0,05

jhonyy9 (jhonyy9):

in the first paranthese only x is over 10 ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

than yes you can write 10(10-x)/10 what equal 10-x

jhonyy9 (jhonyy9):

yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

ok

OpenStudy (anonymous):

and the second term -(10-x)= 10+x

jhonyy9 (jhonyy9):

10-x-(10-x) =10-x-10+x=0 yes ?

OpenStudy (anonymous):

why =0, what about the 0.05?

jhonyy9 (jhonyy9):

now we continue just step by step ok ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please continue

jhonyy9 (jhonyy9):

so there 10-x and -(10-x) yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so this how many will be ?

jhonyy9 (jhonyy9):

comen -(10-x) = ???

OpenStudy (anonymous):

-10+x

jhonyy9 (jhonyy9):

you dont like to solve it ?

jhonyy9 (jhonyy9):

this is very easy

OpenStudy (anonymous):

solve what 10-x-10+x?

jhonyy9 (jhonyy9):

yes first

jhonyy9 (jhonyy9):

and we continue

jhonyy9 (jhonyy9):

what is

jhonyy9 (jhonyy9):

result

OpenStudy (anonymous):

well they cancel so they =0

jhonyy9 (jhonyy9):

ok so till -1/100 there are 0 yes ?

OpenStudy (anonymous):

yes and -1/100(10-x)= -1/10+x

OpenStudy (anonymous):

and =0.05 and be converted into 1/20

jhonyy9 (jhonyy9):

moment please first this (10-x) is in denominator or is over 100 ?

OpenStudy (anonymous):

I am not sure.

jhonyy9 (jhonyy9):

h

jhonyy9 (jhonyy9):

how is in your book ?

OpenStudy (anonymous):

like I have written it out

OpenStudy (anonymous):

I thought it was distributive property

jhonyy9 (jhonyy9):

hence -1 --- (10-x) 100

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so than is hence there we have -(10-x)/100 = 5/100 is it ok ?

jhonyy9 (jhonyy9):

sorry i dont undastand it ?

jhonyy9 (jhonyy9):

comen

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

this is imadiatly ended

jhonyy9 (jhonyy9):

yes ?

OpenStudy (anonymous):

I am kinda lost a little

OpenStudy (anonymous):

(10-x)/100 = 5/100, where is the -x??

jhonyy9 (jhonyy9):

look on what i have wrote in last equation

OpenStudy (anonymous):

is 5/100 the 0.05?

jhonyy9 (jhonyy9):

why not is equal ?

OpenStudy (anonymous):

can that not be reduced to 1/20?

jhonyy9 (jhonyy9):

this is possibille but if you check it in the link term we have like denominator 100

jhonyy9 (jhonyy9):

yes ?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think I can solve it now

jhonyy9 (jhonyy9):

because there is -(10-x)/100 for one easy step is if we us 5/100 there too

OpenStudy (anonymous):

would give give me 1 minute

jhonyy9 (jhonyy9):

hope that you have understood it all now ,yes ?

OpenStudy (anonymous):

I see my mistakes yes

OpenStudy (anonymous):

I have hit a wall

OpenStudy (anonymous):

shouldnt it be 10+x/100?

jhonyy9 (jhonyy9):

yes this not is 10 +x

jhonyy9 (jhonyy9):

this is ???

jhonyy9 (jhonyy9):

comen this is the last step

jhonyy9 (jhonyy9):

courage

OpenStudy (anonymous):

what do I do with the remaining X?

jhonyy9 (jhonyy9):

there is -(10-x)/10=5/100 yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

sorry -(10-x)/100=5/100

jhonyy9 (jhonyy9):

yes ?

jhonyy9 (jhonyy9):

so how you know if the denominators are the same

OpenStudy (anonymous):

because their both 100, but why is it -x and not +x?

jhonyy9 (jhonyy9):

what we have there -(10-x) yes ?

OpenStudy (anonymous):

The term is -1 --- (10-x) 100 does -1*-x=+x

jhonyy9 (jhonyy9):

so how you know it ? -(-x)=???

OpenStudy (anonymous):

because dont you distribute the -1?

jhonyy9 (jhonyy9):

how many is minus multiply minus ?

OpenStudy (anonymous):

I do not understand that

jhonyy9 (jhonyy9):

why ? there we have -(10-x)/100 =5/100 because the denominators are the same we can write -(10-x)=5

jhonyy9 (jhonyy9):

-(10-x) is difficile for you ?

jhonyy9 (jhonyy9):

why ?

jhonyy9 (jhonyy9):

ther -1 multipli

jhonyy9 (jhonyy9):

with 10

jhonyy9 (jhonyy9):

and -1 mult

jhonyy9 (jhonyy9):

ply with -x

OpenStudy (anonymous):

because doesnt a negative * negative = positive?

jhonyy9 (jhonyy9):

but is

jhonyy9 (jhonyy9):

- * - =+

jhonyy9 (jhonyy9):

o

jhonyy9 (jhonyy9):

or you need to learn again

OpenStudy (anonymous):

maybe, may be go back please to -1/100(10-x)